OpenStudy (anonymous):

Sophie drove 225 mi south through the Smoky Mountains. Snow during her return trip made her average speed on the return trip 30 mph slower. The total driving time was 8 hr. Find Sophie’s average sped on each part of the trip.

OpenStudy (anonymous):

@Nnesha

jimthompson5910 (jim_thompson5910):

let x = speed on her first part of the trip x-30 is the speed on the second part of her trip coming back

jimthompson5910 (jim_thompson5910):

d = r*t 225 = x*t t = 225/x is the time it takes to drive 225 mi at x mph d = r*t 225 = (x-30)*t t = 225/(x-30) is the time it takes to drive 225 mi at x-30 mph combine the two times and set that equal to 8 \[\Large \frac{225}{x}+\frac{225}{x-30} = 8\] now solve for x